# Switching the Loss Reduces the Cost in Offline Reinforcement Learning

### v2

@@ Abstract
We propose training fitted Q-iteration with log-loss (FQI-log) for offline reinforcement learning (RL). We show that the number of samples needed to learn a near-optimal policy with FQI-log scales with the accumulated cost of the optimal policy, which is, e.g., small or even zero in goal-oriented environments where acting optimally reliably achieves the goal. As a byproduct, we provide a general framework for proving small-cost bounds in offline RL. Moreover, we empirically verify that FQI-log uses fewer samples than FQI trained with squared loss on problems where the optimal policy reliably achieves the goal.
@@ 1 Introduction
In offline reinforcement learning (RL), also known as batch RL, we often want agents that learn how to achieve a goal from a fixed dataset using as few samples as possible. A standard approach in this setting is fitted Q-iteration (FQI) (Ernst et al., 2005), which iteratively minimizes the regression error on the batch dataset. In this work we propose a simple and principled improvement to FQI, using log-loss (FQI-log), and prove that it can achieve a much faster convergence rate. In particular, the number of samples it requires to learn a near-optimal policy scales with the cost of the optimal policy, leading to a so-called small-cost bound, the RL analogue of a small-loss bound in supervised learning. We highlight that FQI-log is the first computationally efficient batch RL algorithm to achieve a small-cost bound.
The intuition behind training FQI with log-loss comes from the observation that not all state-action-to-next-state transitions are equally noisy. Whereas training with squared loss amounts to assuming the noise is homoscedastic across transitions, training with log-loss learns to emphasize less noisy transitions and de-emphasize more noisy transitions. Since in RL typically some transitions are noisier than others, learning to emphasize less noisy transitions leads to more sample efficient learning.
Our paper is structured as follows. Section 2 contains a description of the problem setting and sets out our notation. In Section 3 we propose FQI-log and motivate the switch to log-loss from squared loss, which is the standard loss for FQI (Riedmiller, 2005; Chen and Jiang, 2019). In Section 4, we present a general framework for deriving small-cost bounds in batch RL, and then use this analysis to prove that FQI-log enjoys a small-cost guarantee. Section 5 presents empirical verification that FQI-log, indeed, requires fewer samples and more reliably achieves goals, when compared to training FQI with squared loss (FQI-sq).
[…]
The main contributions of this work can be summarized as:
  1. 1.
    We propose training FQI with log-loss and prove it enjoys a small-cost bound. This is the first efficient batch RL algorithm that achieves a small-cost bound.
  2. 2.
    We present new analysis showing that the Bellman optimality operator is a contraction with respect to the Hellinger distance.
  3. 3.
    We empirically evaluate FQI-log and FQI-sq on Mountain Car (Moore, 1990) and Atari 2600 games (Bellemare et al., 2013), demonstrating that FQI-log consistently outperforms FQI-sq.
@@
We also highlight an intermediate result that may be of independent interest. Specifically, we present a general result that decomposes the suboptimality gap of the value of a greedy policy induced by some value function, $f$, into the product of a small-cost term and the pointwise triangular deviation of $f$ from $q^{\star}$, the value function of the optimal policy.
The main limitation of our approach is that FQI-log requires a more stringent boundedness assumption than FQI-sq. We make Assumption 4.6 because log-loss cannot handle unbounded inputs. In comparison, squared loss can handle unbounded inputs and thus FQI-sq only requires that the function class and intermediate costs (or rewards) map to bounded intervals (Chen and Jiang, 2019).
@@ 2 Preliminaries
We consider an infinite-horizon discounted Markov Decision Process (MDP) defined as $M=(\mathcal{S},\mathcal{A},P,c,\gamma,\eta_{1})$, where $\mathcal{S}$ is the (finite, but arbitrarily large) state space, $\mathcal{A}$ is the (finite) action space, $P:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{M}_{1}(\mathcal{S})$ is the transition function ($\mathcal{M}_{1}(\cdot)$ is the probability simplex), $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the cost function, $\gamma\in[0,1)$ is the discount factor for future costs, and $\eta_{1}\in\mathcal{M}_{1}(\mathcal{S})$ is the distribution from which states are drawn when the process is initialized. size=,color=red!20!white,]Cs: Is $\eta_{1}$ part of the MDP spec? size=,color=green!20!white,]AA: I’ve seen it included, like in our VTR work. Though maybe it is a property of $\mathcal{S}$
Remark 2.1.
We assume that $\mathcal{S}$ is finite solely for exposition. This allows us to simplify the presentation of our analysis and focus on the most salient details of the proof, avoiding the cumbersome measure-theoretic notation required to reason about infinite sets.
A stationary (stochastic) policy $\pi:\mathcal{S}\rightarrow\mathcal{M}_{1}(\mathcal{A})$ is a mapping from states to distributions over actions. A nonstationary policy $\tilde{\pi}_{t}=\{\tilde{\pi}_{h}\}_{h=1}^{t}$ is a set of $t$ functions where $\tilde{\pi}_{h}:(\mathcal{S}\times\mathcal{A})^{h-1}\times\mathcal{S}\rightarrow\mathcal{M}_{1}(\mathcal{A})$ maps a length $h-1$ state-action sequence and a state to a distribution over actions. size=,color=orange!20!white,]SR: Should $t$ be a subscript here? This conflicts with itself. Fixing the start state $s$, policy $\pi$ induces a distribution $\mathbb{P}_{\pi,s}$ over trajectories $S_{1},A_{1},C_{1},S_{2},A_{2},C_{2},\dots,$ where $S_{1}=s$, $A_{1}\sim\pi(S_{1})$, $C_{1}=c(S_{1},A_{1})$, $S_{2}\sim P(S_{1},A_{1})$, etc. The expected total discounted cost over trajectories starting in $s$ quantifies the policy’s performance when initialized in state $s$. We collect these expectations in the state-value function of $\pi$, $v^{\pi}:\mathcal{S}\to\mathbb{R}$, which is defined by $v^{\pi}(s)=\mathbb{E}_{\pi,s}[\sum_{h=1}^{\infty}\gamma^{h-1}C_{h}]$, where $\mathbb{E}_{\pi,s}$ is the expectation operator corresponding to $\mathbb{P}_{\pi,s}$. We also denote the expected value of $v^{\pi}$ with respect to the initial state distribution $\eta_{1}$ by $\bar{v}^{\pi}=\left\langle\eta_{1},\,v^{\pi}\right\rangle$, where for vectors $a,b$ of matching dimension $\left\langle a,\,b\right\rangle=\sum_{i}a_{i}b_{i}$ denotes the usual inner product. The action-value function underlying $\pi$, $q^{\pi}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, is defined as
@@
for $(s,a)\in\mathcal{S}\times\mathcal{A}$. size=,color=orange!20!white,]SR: Should we replace the sum by $P(s,a),v^{\pi}\right>$? The state- and action-value functions are related by the identity ${v^{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)q^{\pi}(s,a)}$. size=,color=red!20!white,]Cs: This is a bad abuse of notation: The reader reading $v^{\pi}$ does not know whether this is a real, or a function. size=,color=orange!20!white,]SR: Is this okay if we write $\bar{v}^{\pi}=\left\langle\eta_{1},\,v^{\pi}\right\rangle\in\mathbb{R}$? I think since $\eta_{1}$ and $v^{\pi}$ are both functions over $\mathcal{S}$ above it is not ambiguous though?
The optimal policy is defined as any policy $\pi^{\star}$ that satisfies $v^{\pi^{\star}}(s)=\min_{\pi}v^{\pi}(s)$ simultaneously for all $s\in\mathcal{S}$. We define the optimal state-value function as $v^{\star}=v^{\pi^{\star}}$ and the optimal action-value function as $q^{\star}=q^{\pi^{\star}}$; due to the finite state and action spaces, both $v^{\star}$ and $q^{\star}$ are well-defined. Any policy that is greedy with respect to $q^{\star}$, i.e. at state $s$ selects only actions $a$ that minimize $q^{\star}(s,a)$, is guaranteed to be optimal. Furthermore, the optimal action-value function $q^{\star}$ satisfies the Bellman equation $q^{\star}=\mathcal{T}q^{\star}$, where $\mathcal{T}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\rightarrow\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ is the Bellman optimality operator,
@@
Finally, we use $\pi_{f}$ to denote a greedy policy induced by $f$, $\pi_{f}(s)\in\mathcal{M}_{1}(\mathop{\rm arg\,min}_{a\in\mathcal{A}}f(s,a))$. When there are multiple such policies, choose one in an arbitrary (systematic) manner to make $\pi_{f}$ well-defined. size=,color=orange!20!white,]SR: Subscripting $\pi$ does conflict with the nonstationary policies being indexed by timesteps. size=,color=orange!20!white,]SR: The old definition was $\pi_{f}(s)=\mathop{\rm arg\,min}_{a\in\mathcal{A}}f(s,a)$, I changed it to make it more precise but maybe it is less clear now.
Batch RL
We assume the agent knows the state space $\mathcal{S}$ and the action space $\mathcal{A}$ but does not know the cost function $c$ nor the transition function $P$. The agent is given a random dataset $D=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ and a class of candidate value functions $\mathcal{F}\subset\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$. The dataset $D$ is constructed by sampling state-action pairs i.i.d. from an unknown exploratory distribution $\mu$: for each $i\in\{1,\dotsc,n\}$, $(S_{i},A_{i})\sim\mu$, $C_{i}=c(S_{i},A_{i})$, and $S_{i}^{\prime}\sim P(S_{i},A_{i})$. Given $D$ and $\mathcal{F}$ the agent’s goal is to find a near-optimal policy.
2.1 Additional Notation
@@ 3 FQI-log: Fitted Q-Iteration with log-loss
The proposed algorithm, FQI-log, which is described in Algorithm 1, is based on the fitted Q-iteration algorithm (FQI) (Ernst et al., 2005; Riedmiller, 2005; Antos et al., 2007). Given a batch dataset, FQI iteratively produces a sequence of $k$ estimates $f_{1},\ldots,f_{k}$ of the action-value function $q^{\star}$. At iteration $j\in[k]$, the algorithm computes $f_{j}$ by minimizing the empirical loss using targets estimated with $f_{j-1}$, the result of the previous iteration. The targets are constructed such that the regression function for a fixed estimate $f\in\mathcal{F}$ is $\mathcal{T}f$. The main difference between the proposed method and the most common variant of FQI is our use of log-loss,
$\ell_{\log}(y;\,\tilde{y})=\tilde{y}\log\frac{1}{y}+(1-\tilde{y})\log\frac{1}{1-y}\,\,(y\in[0,1],\tilde{y}\in\mathbb{R})\,,$ (2)
to measure the deviation between the targets and the estimates. This loss is well-defined only when $y\in[0,1]$. In our algorithm the first argument of $\ell_{\log}$ is a predicted value $f(S_{i},A_{i})$ with $f\in\mathcal{F}$. Thus running FQI-log requires the range of all functions in $\mathcal{F}$ to lie in $[0,1]$. size=,color=red!20!white,]Cs: it is unclear to me whether we need a restriction on $\tilde{y}$ other than it is nonnegative perhaps. We need to check the analysis. In Dylan’s paper, I guess.. size=,color=green!20!white,]AA: Dylan has that the predictor $y$ is in $[0,1]$. Cs: But does their analysis use this? size=,color=orange!20!white,]SR: If $\tilde{y}$ is outside of $[0,1]$ then the loss could become negative. We also need $\tilde{y}$ to have nonnegative expectation for the expectation to be the minimizer in the next paragraph, and I think the expectation should be $\leq 1$.
Note that previous work on FQI employed the squared loss $\ell_{\text{sq}}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, defined as $\ell_{\text{sq}}(y;\,\tilde{y})=(y-\tilde{y})^{2}$, instead of the log-loss $\ell_{\log}$. Both squared loss and log-loss have the property that for a variance bounded, nonnegative random variable $X$, $\mathbb{E}\,[X]=\mathop{\rm arg\,min}_{m\in\mathbb{R}}\mathbb{E}\,[\ell(m,X)]$ where $\ell$ is either $\ell_{\log}$ or $\ell_{\text{sq}}$. size=,color=red!20!white,]Cs: Indeed, if the variance of $X$ is bounded, $\mathbb{E}\,{\ell_{\log}(m,X)}=f(\mathbb{E}\,{X})$ is finite. In fact, only the mean needs to be bounded for that log-loss. size=,color=orange!20!white,]SR: In my simple (naive?) derivation that this is the minimizer for log-loss, for the function to be convex it’s required that $\mathbb{E}\,{\tilde{y}}\in[0,1]$, is this not necessary generally?
The differences between log-loss and squared loss are made apparent in regression. As is well known, the ideal loss in regression compensates when the outcomes of the observed data points have unequal variances, a condition often called “heteroscedastic regression”. In heteroscedastic regression, it is desirable to de-emphasize noisy outcomes, i.e. outcomes whose noise has high variance, and emphasize less noisy outcomes. As opposed to squared loss, which treats each data point as having equal-variance outcomes (“homoscedastic regression”), log-loss assigns a higher cost size=,color=green!20!white,]AA: Very ambiguous? What is cost here. Cs: cost is what you pay. I think cost work better here. to any predictor, $f$, that is closer to zero, which is obvious from its definition in Eq. 2. size=,color=orange!20!white,]SR: Should this say “to any predictor that is closer to 0 or 1”? This makes sense: if the response is nonnegative valued and bounded, predicting that its mean is small implies that the variance of the response is also small, hence, higher cost should be incurred when a small value is predicted.
Another insight related to these losses is that they correspond to maximizing the likelihood of a regression function with fixed covariates and noise distributions. Squared loss is equivalent to minimizing the empirical loss of centered Gaussian distributions with variances equal across the data points. Log-loss corresponds to minimizing the empirical loss of centered Bernoulli distributions, with variances dependent on the Bernoulli parameters. size=,color=red!20!white,]Cs: I purposefully not talk about the boundedness and the “high” edge.
With this intuition in mind, when can we expect the choice of the loss to matter? Clearly, the choice matters when the regression function gives values either close to 00 or close to $1$, either boundary of the allowed range, i.e. $[0,1]$. Since our goal is to minimize cost, which is expressed with the help of these regression functions, our intuition suggests that log-loss may induce better policies when the ultimate target, $q^{\star}$, is small. We will affirm this intuitive reasoning, below, in our main result.
@@
The motivation to switch to log-loss is due to Foster and Krishnamurthy (2021) who studied the problem of learning a near-optimal policy in contextual bandits, both in the batch setting and the online setting. They noticed that switching to log-loss from squared loss allows bounding the suboptimality of the policy found, say in the batch setting after seeing $n$ contexts, via a term that scales with $\sqrt{\bar{v}^{\star}/n}+1/n$. This is an improvement from the usual $\sqrt{1/n}$ bound derived when analyzing squared loss, which is worst-case in nature. For log-loss, a significant speedup to $1/n$-type convergence is achieved when $\bar{v}^{\star}$, the expected cost of using the optimal policy, is small (cf. Section 3.1 of their paper). They complemented the theory with convincing empirical demonstrations. Our results take a similar form. While we reuse some of their results and techniques, our analysis deviates significantly from theirs. In particular, our analysis must be adapted to handle the multistage structure present in RL and to avoid an unnecessary dependence on the actions.
@@
The astute reader may wonder whether switching to log-loss is really necessary for achieving small-cost bounds. As it turns out, the switch is necessary, as attested to by an example constructed in Foster and Krishnamurthy (2021). In this example, squared loss is shown to be unable to take advantage of small optimal costs, in contrast to log-loss. (cf. Theorem 2 of Foster and Krishnamurthy (2021)).

### v5

# Switching the Loss Reduces the Cost in Batch Reinforcement Learning

@@ Abstract
We propose training fitted Q-iteration with log-loss (FQI-log) for batch reinforcement learning (RL). We show that the number of samples needed to learn a near-optimal policy with FQI-log scales with the accumulated cost of the optimal policy, which is zero in problems where acting optimally achieves the goal and incurs no cost. In doing so, we provide a general framework for proving small-cost bounds, i.e. bounds that scale with the optimal achievable cost, in batch RL. Moreover, we empirically verify that FQI-log uses fewer samples than FQI trained with squared loss on problems where the optimal policy reliably achieves the goal.
Machine Learning, ICML
@@ 1 Introduction
In batch reinforcement learning (RL), also known as offline RL, the goal is to learn a good policy from a fixed dataset. A standard approach in this setting is fitted Q-iteration (FQI) (Ernst et al., 2005), which iteratively obtains a sequence of value functions by fitting the next value function to a target that is based on the data and the previously obtained value function. In this work we propose a simple and principled improvement to FQI, using log-loss (FQI-log), which is applicable when the returns along trajectories lie in a bounded interval. We prove that the number of samples the new method requires to learn a near-optimal policy scales with the cost of the optimal policy, leading to a so-called small-cost bound, the RL analogue of a small-loss bound in supervised learning. Such bounds predict improved sample efficiency in goal oriented RL tasks where the goal is reliably achievable and the cost is set up to penalize failure in achieving the goal; a prediction we validate empirically. We highlight that FQI-log is the first computationally efficient batch RL algorithm to achieve a small-cost bound, provided that a regression oracle is available; a condition that can be met, e.g., when logit models are used in FQI-log.
Most earlier works in batch RL focused on algorithms that achieve the optimal worse-case dependence on the number of samples required to learn a near-optimal policy (Munos, 2003; Antos et al., 2007; Chen & Jiang, 2019; Xie & Jiang, 2021). The only work prior to ours that is known to adaptively improve sample efficiency when facing a task with near-zero optimal cost is due to Wang et al. (2023), who obtain small-cost bounds for finite-horizon batch RL problems but using the so-called “distributional RL” approach. In this approach, one solves the regression problems arising when estimating the distribution of a policy’s accumlated cost. The inspiration for our work comes from this work, combined with the insights of Abeille et al. (2021); Foster & Krishnamurthy (2021) who showed that, in the simpler bandit problems, log-loss alone is sufficient for obtaining small-cost bounds.
Why log-loss gives a small-cost advantage is subtle. When used with an unrestricted model class (think: finite state-action space, “tabular learning”), both log- and squared losses achieve small cost bounds because they share the same minimizers, which predict the empirical mean of responses for all inputs. However, when the model class is restricted (the only practical case for large problems), log-loss and squared loss will trade off errors at the various inputs differently. Specifically, with log-loss, the penalty for predicting a value far from an observed mean increases rapidly as the observed mean gets close to the boundary of its range, an effect that is absent with squared loss. Consequently, log-loss will favor predictors that fit well to those observed values that are near the boundary of the range, making the learning process disregard large variance observed values, stabilizing learning, a property not shared when a squared loss is used. As a result, as we show, under suitable additional technical assumptions, log-loss based FQI achieves small-cost bounds, a property that is not shared when squared loss is used with FQI.
@@
The main contributions of this work can be summarized as: (i) We propose training FQI with log-loss and prove it enjoys a small-cost bound. This is the first efficient batch RL algorithm that achieves a small-cost bound. When showing this result, we make two technical contributions that may be of independent interest: (ii) We show that the Bellman optimality operator is a contraction with respect to the Hellinger distance, a result; (iii) We present a general result that decomposes the suboptimality gap of the value of a greedy policy induced by some value function, $f$, into the product of a small-cost term and the pointwise triangular deviation of $f$ from $q^{\star}$, the value function of the optimal policy.
@@ 2 Preliminaries
In this section we review some definitions and concepts of Markov Decision Processes (MDPs) and define the notation used. Readers unfamiliar with the basic theory of MDPs are recommended to consult the book of Bertsekas (2019), or that of Szepesvári (2010). All results mentioned in this section can be found in these works.
An infinite-horizon discounted Markov Decision Process (MDP) is given by a tuple size=,color=orange!20!white,]SR: we already introduced the acronym just above, is this desired? $M=(\mathcal{S},\mathcal{A},P,c,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P$ is a transition function, $c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a cost function and $\gamma\in[0,1)$ is a discount factor. We only consider MDPs with finite action spaces. Furthermore, for simplicity, we assume that the state space is finite. Among other things, [^1] We assume that the state space $\mathcal{S}$ is finite solely for exposition. This allows us to simplify the presentation of our analysis and focus on the most salient details of the proof, avoiding the cumbersome measure-theoretic notation required to reason about infinite sets. this allows writing the transition function as $P:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{M}_{1}(\mathcal{S})$, where $\mathcal{M}_{1}(\mathcal{S})$ denotes the set of probability distributions over $\mathcal{S}$. Since the set $\mathcal{S}$ is finite, any element of $\mathcal{M}_{1}(\mathcal{S})$ can and will be identified with its probability mass function. The notation $\mathcal{M}_{1}(\mathcal{X})$ will be used in the same way to denote the set of probability distributions over an arbitrary finite set $\mathcal{X}$ and we will perform the same identification.
A (general) policy $\pi=(\pi_{h})_{h=1}^{\infty}$ is a sequence of functions $\pi_{h}:(\mathcal{S}\times\mathcal{A})^{h-1}\times\mathcal{S}\rightarrow\mathcal{M}_{1}(\mathcal{A})$. Fixing the start state $s$, a policy $\pi$ induces a distribution $\mathbb{P}_{\pi,s}$ over trajectories $S_{1},A_{1},C_{1},S_{2},A_{2},C_{2},\dots,$ where $S_{1}=s$, $A_{1}\sim\pi_{1}(S_{1})$, $C_{1}=c(S_{1},A_{1})$, $S_{2}\sim P(S_{1},A_{1})$, $A_{2}\sim\pi_{2}(S_{1},A_{1},S_{2})$: the policy is used to govern the selection of actions, while the transition dynamics of the MDP governs the evolution of the states in response to the chosen actions. We will also need stationary Markov policies, where the choice of the action in any timestep $h$ only depends on the last state visited. Thus, such policies can be identified with a map $\pi:\mathcal{S}\rightarrow\mathcal{M}_{1}(\mathcal{A})$, an identification which we will employ in what follows.
The expected total discounted cost over trajectories starting in $s$ quantifies the policy’s performance when initialized in state $s$. We collect these expectations in the state-value function of $\pi$, $v^{\pi}:\mathcal{S}\to\mathbb{R}$, which is defined by $v^{\pi}(s)=\mathbb{E}_{\pi,s}[\sum_{h=1}^{\infty}\gamma^{h-1}C_{h}]$, where $\mathbb{E}_{\pi,s}$ is the expectation operator corresponding to $\mathbb{P}_{\pi,s}$. For convenience, throughout this paper we assume that costs are normalized so that the sum of discounted costs along any trajectory satisfies
$0\leq\sum_{h=1}^{\infty}\gamma^{h-1}C_{h}\leq 1\,.$ (1)
size=,color=red!20!white,]Cs: We may need to highlight this better.By appropriately rescaling the costs, this constraint can always be satisfied (when the state space is not finite, one needs that the above infinite sums are uniformly bounded to be able to do this). size=,color=red!20!white,]Cs: comment after the result on how rescaling and shifting impacts the bounds.
The action-value function of $\pi$, $q^{\pi}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, is defined as
@@
where $(s,a)\in\mathcal{S}\times\mathcal{A}$ and, by abusing notation, we use $P(s^{\prime}|s,a)$ to denote the probability of landing in state $s^{\prime}$ when action $a$ is taken in state $s$. For a stationary Markov policy $\pi$, the state- and action-value functions are related by the identity $v^{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)q^{\pi}(s,a)$. Here, and in what follows, abusing notation once again, $\pi(a|s)$ denotes the probability that is assigned by $\pi(s)$ to action $a\in\mathcal{A}$.
The optimal policy is defined as any policy $\pi^{\star}$ that satisfies $v^{\pi^{\star}}(s)=\min_{\pi}v^{\pi}(s)$ simultaneously for all $s\in\mathcal{S}$. Such a policy exists in our case. We define the optimal state-value function as $v^{\star}=v^{\pi^{\star}}$ and the optimal action-value function as $q^{\star}=q^{\pi^{\star}}$. Any policy that is greedy with respect to $q^{\star}$, i.e., at state $s$ selects only actions $a$ that minimize $q^{\star}(s,\cdot)$, is guaranteed to be optimal. Furthermore, the optimal action-value function $q^{\star}$ satisfies the Bellman optimality equation $q^{\star}=\mathcal{T}q^{\star}$, where $\mathcal{T}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\rightarrow\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ is the Bellman optimality operator, defined via
@@
Finally, we use $\pi_{f}$ to denote a greedy policy induced by $f$: $\pi_{f}(s)=\mathop{\rm arg\,min}_{a\in\mathcal{A}}f(s,a)$. When there are multiple such policies, we choose one in an arbitrary (systematic) manner to make $\pi_{f}$ well-defined. size=,color=orange!20!white,]SR: Subscripting $\pi$ does conflict with the nonstationary policies being indexed by timesteps. Cs: Different types; not a problem. size=,color=orange!20!white,]SR: The old definition was $\pi_{f}(s)=\mathop{\rm arg\,min}_{a\in\mathcal{A}}f(s,a)$, I changed it to make it more precise but maybe it is less clear now. Cs: I don’t see the difference.
3 Problem Definition: Batch RL
[…]
The data is assumed to satisfy the following assumption:
Assumption 3.1 (Data/MDP properties).
We have that $D_{n}=\{(S_{i},A_{i},C_{i},S_{i}^{\prime})\}_{i=1}^{n}$ is a sequence of independent, identically distributed random variables such that $(S_{i},A_{i})\sim\mu$, $C_{i}=c(S_{i},A_{i})$ and $S_{i}^{\prime}\sim P(S_{i},A_{i})$. Furthermore, Equation 1 holds and $c$ is a nonnegative function.
In this assumption the constraints on $C_{i}$ and $S_{i}^{\prime}$ will help the learner to discover some information about the costs and the transition structure of the MDP. The independence assumption is made to simplify the analysis. size=,color=red!20!white,]Cs: Do we know what happens if we remove this..? The distribution $\mu$ will be further restricted to be “sufficiently exploratory”. To state this assumption, the following definition will be useful:
Definition 3.1 (Admissible distribution).
Note that what constitutes an admissible distribution depends on $\eta_{1}$. With this, the assumption that constrains $\mu$ is as follows:
Assumption 3.2 (Finite concentrability coefficient).
There exists $C<\infty$ such that for all admissible distributions $\nu$ of $M$, it holds that
$\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{\nu(s,a)}{\mu(s,a)}\leq C\,.$ (3)
Note that if $\mu$ is positive over $\mathcal{S}\times\mathcal{A}$, the assumption is satisfied. By only considering admissible distributions, we allow $\mu$ to be “concentrated” on states that are “relevant” in the sense that they are visited by some policy in some time step with a large probability when starting from $\eta_{1}$.
In addition to having access to the data, we will also assume that the learner is given access to a set of functions, $\mathcal{F}$ that map state-action pairs to reals: $\mathcal{F}\subset\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$. Ideally, the set $\mathcal{F}$ allows the learner to reason about the optimal action-value function. To make this possible, the following assumptions are made on $\mathcal{F}$:
Assumption 3.3 (Realizability).
$q^{\star}\in\mathcal{F}$.
Assumption 3.4 (Completeness).
For all $f\in\mathcal{F}$ we have that $\mathcal{T}f\in\mathcal{F}$.
Assumptions 3.1, 3.2, 3.3 and 3.4 are commonly made, in various forms, when analyzing fitted Q-iteration (Farahmand, 2011; Pires & Szepesvári, 2012; Chen & Jiang, 2019). Assumption 3.2 ensures that all admissible distributions are covered by the exploratory distribution $\mu$, i.e., that “$\mu$ is sufficiently exploratory”. Assumption 3.3 (“realizability”) guarantees that the optimal action-value function, our ultimate target, lies in our function class. Assumption 3.4 states that the function class $\mathcal{F}$ is closed under the Bellman optimality operator $\mathcal{T}$. When $\mathcal{F}$ is closed, completeness is easily seen to imply realizibility (see also footnote 10 of Chen & Jiang (2019)). Note that Assumption 3.4 is necessary, this is due to a result by Foster et al. (2021) which states that assuming both a finite concentrability coefficient and a realizable function class are not sufficient for sample efficient batch value function approximation. For a more detailed discussion of the last three assumptions, we refer the reader to Sections 4 and 5 of Chen & Jiang (2019).
Research question
As is well known, under the above assumptions, and assuming that a regression oracle is available to find the empirical minimizer of regression problems defined over $\mathcal{F}$ with the squared loss, the so-called fitted Q-iteration (FQI) algorithm (Ernst et al., 2005; Riedmiller, 2005; Antos et al., 2007) produces a policy such that with high probability $\bar{v}^{\pi}\leq\bar{v}^{\star}+\tilde{O}(\sqrt{CN/n})$, where $N$ is a measure that characterizes the “richness” of $\mathcal{F}$ and $\tilde{O}$ hides logarithmic factors (Antos et al., 2007). The main question investigated in this paper is whether this result can be improved to $\bar{v}^{\pi}\leq\bar{v}^{\star}+\tilde{O}(\sqrt{CN\bar{v}^{\star}/n})+O(1/n)$. The significance of such small cost results is that the same data can produce a significantly better policy when $\bar{v}^{\star}$ is near zero (note that $\bar{v}^{\star}\geq 0$). Alternatively, the number of samples required to achieve a given level of suboptimality can be significantly smaller if an algorithm satisfies a small-cost bound.
Additional notation
@@ 4 FQI-log: Fitted Q-Iteration with log-loss
The proposed algorithm, FQI-log, which is described in Algorithm 1, is based on the earlier mentioned fitted Q-iteration algorithm (FQI) (Ernst et al., 2005; Riedmiller, 2005; Antos et al., 2007). Given a batch dataset, FQI iteratively produces a sequence of $k$ approximations $f_{1},\ldots,f_{k}$ to the action-value function $q^{\star}$. At iteration $j\in[k]$, the algorithm computes $f_{j}$ by minimizing the empirical loss using targets computed with the help of $f_{j-1}$, the estimate produced in the previous iteration. The targets are constructed such that the regression function for a fixed estimate $f\in\mathcal{F}$ is $\mathcal{T}f$. The main difference between the proposed method and the most common variant of FQI is our use of log-loss,
$\ell_{\log}(y,\tilde{y})=\tilde{y}\log\frac{1}{y}+(1-\tilde{y})\log\frac{1}{1-y}\,,$ (4)
to measure the deviation between a prediction ($y\in[0,1]$) and a target $(\tilde{y}\in[0,1]$), where to allow $y,\tilde{y}\in\{0,1\}$, we use $00\cdot\log 0=$ 0. The restriction that $\tilde{y}\in[0,1]$ means that $\ell_{\log}(\cdot,\tilde{y})$ is convex. In our algorithm the first argument of $\ell_{\log}$ is a predicted value $f(S_{i},A_{i})$ with $f\in\mathcal{F}$. Since this needs to also belong to $[0,1]$, running FQI-log requires the range of all functions in $\mathcal{F}$ to lie in $[0,1]$. size=,color=red!20!white,]Cs: it is unclear to me whether we need a restriction on $\tilde{y}$; per the next paragraph we need that whatever random response goes in place of $\tilde{y}$ has an expected value in $[0,1]$. But do we need it is in $[0,1]$? Maybe convexity is used somewhere in the analysis?
Note that previous work on FQI employed the squared loss $\ell_{\text{sq}}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, defined as $\ell_{\text{sq}}(y,\tilde{y})=(y-\tilde{y})^{2}$, instead of the log-loss $\ell_{\log}$. Both squared loss and log-loss have the property that for a random variable $X$ taking values in $[0,1]$ with probability one, $\mathbb{E}\,[X]=\mathop{\rm arg\,min}_{m\in\mathbb{R}}\mathbb{E}\,[\ell(m,X)]$, where $\ell$ is either $\ell_{\log}$ or $\ell_{\text{sq}}$. In particular, if $\mathcal{F}=[0,1]^{\mathcal{S}\times\mathcal{A}}$, both log-loss and squared loss will give rise to the same sequence $f_{1},\dots,f_{k}$.
The differences between log-loss and squared loss are made apparent when $\mathcal{F}\neq[0,1]^{\mathcal{S}\times\mathcal{A}}$. In this case, when $\tilde{y}$ is near 00 or $1$, $\ell_{\log}(\cdot,\tilde{y})$ increases rapidly as $y$ deviates from $\tilde{y}$. As such, when errors need to be traded off at different inputs, log-loss will end up favoring predictors that predict values closer to observed targets when the targets are near 00 (or $1$), and will put less weight on observed targets in the middle of the $[0,1]$ range. When a true value lies near 00 (or $1$), the observed value (bound to the range $[0,1]$) must be closer to the true value, which means that the observed value is also close to 00 (or $1$). While it may happen that an observed value is close to 00 (or $1$) while its mean is far from it, this is rare: For this to happen, the observed value has to have a large variance. As such, favoring to predict observed values near the 00 or $1$ as opposed to paying equal attention to all datapoints (which is what the squared loss based predictors do) is beneficial, and, in particular it pays off when some of the true targets are near 00 (or $1$): The situation that arises when the optimal cost is near zero.
@@
The motivation to switch to log-loss is due to Foster & Krishnamurthy (2021) who studied the problem of learning a near-optimal policy in contextual bandits, both in the batch and the online settings. They noticed that switching to log-loss from squared loss allows bounding the suboptimality of the policy found, say in the batch setting after seeing $n$ contexts, via a term that scales with $\sqrt{\bar{v}^{\star}/n}+1/n$. This is an improvement from the usual $\sqrt{1/n}$ bound derived when analyzing squared loss, which is worst-case in nature. For log-loss, a significant speedup to $1/n$-type convergence is achieved when $\bar{v}^{\star}$, the expected cost of using the optimal policy, is small (cf. Section 3.1 of their paper). They complemented the theory with convincing empirical demonstrations. Our results take a similar form. While we reuse some of their results and techniques, our analysis deviates significantly from theirs. In particular, our analysis must be adapted to handle the multistage structure present in RL and to avoid an unnecessary dependence on the actions.
@@
The astute reader may wonder whether switching to log-loss is really necessary for achieving small-cost bounds. As it turns out, the switch is necessary, as attested to by an example constructed by Foster & Krishnamurthy (2021). In this example, in contrast to log-loss, squared loss is shown to be unable to take advantage of small optimal costs (cf. Theorem 2 of Foster & Krishnamurthy (2021)).